OpenStudy (anonymous):

Find a piecewise definition of f that does not involve the absolute value function. --->

OpenStudy (anonymous):

\[f(x)=\left| x+4 \right|\]

OpenStudy (anonymous):

square it and take square root

OpenStudy (anonymous):

\[\sqrt{(x+4)^2}\]

OpenStudy (anonymous):

i think we are over complicating it...they actually want a piecewise function this time.

OpenStudy (anonymous):

I can do this nicely in LaTeX, one second

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

\[f(x) = |x+4| = \left\{\begin{array}{c|c} x + 4 & x \geq -4 \\ - x - 4 & x < -4 \end{array} \right.\]

OpenStudy (anonymous):

yeah thats what they were looking for. we can one-up them though with our: \[\sqrt{(x+4)^2}\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[ f(x) = \left\{ \begin{array}{lr} -x-4 & : x <-4\\ x+4 & : x \geq 4 \end{array} \right.\]

OpenStudy (zarkon):

I prefer commas \[f(x) = |x+4| = \left\{\begin{array}{rc} x + 4, & x \geq -4 \\ - x - 4, & x < -4 \end{array} \right.\] and right justifying the terms :)

OpenStudy (anonymous):

a line is nice and clean

OpenStudy (anonymous):

yeah yours is better

OpenStudy (anonymous):

\[f(x) = \left\{ \begin{array}{lr} -x-4 & : x <-4\\ x+4 & : x \geq -4 \end{array} \right.\]

OpenStudy (zarkon):

or \[f(x) = |x+4| = \left\{\begin{array}{rcc} x + 4 & \text{if} & x \geq -4 \\ - x - 4& \text{if} & x < -4 \end{array} \right.\]

OpenStudy (anonymous):

Yes I've seen that format often too

OpenStudy (anonymous):

show off!